18. We are evaluating a project that costs $1.675,000, has a six-year life, and has no salvage value. Assume that depreciation is straight-line to zero over the life of the project. Sales are projected at 91,000 units per year. Price per unit is 535.95, variable cost per unit is $21.40, and fixed costs are $775,000 per year. The tax rate is 35 percent, and we require a return of 11 percent on this project Required Suppose the projections given for price, quantity, variable costs, and fixed costs are all accurate to within ±10 percent. Calculate the best-case and worst-case NPV figures. (Do not round intermediate calculations. Negative amounts should be indicated by a minus sign. Round your answers to 2 decimal places (eg. 32.16).) NPV Best-case Worst-case
Answer #1

For the best case scenario, the price and quantity increase by 10%. and the variable costs and fixed costs both decrease by 10%.

For the worst case scenario, the price and quantity decrease by 10%. and the variable costs and fixed costs both increase by 10%.

We will use the tax shield approach to calculate the OCF. The OCF is:

OCFbase = [(P – VC)Q – FC](1 –t) + Depreciation(t)

OCFbase = [($35.95 – 21.40)(91 000) – $775 000](0.65) + 0.35($1 675 000/6) = $356882.5 + 97708.33

OCF base = $ 454590.83

Now we can calculate the NPV using our base-case projections. There is no salvage value or NWC, so theNPV is:

NPVbase = –$1 675 000 + 454590.83 * PVIFA

=–$1675000 + 454590.83 * 4.230538

= 248163.71

OCFbest case = [(P – VC)Q – FC](1 –t) + Depreciation(t)

OCFbest = [($39.545 – 19.26)(100100) – $697500](0.65) + 0.35($1 675 000/6) = $866468.525 + 97708.33

OCF best = $ 964176.855

Now we can calculate the NPV using our best-case projections. There is no salvage value or NWC, so theNPV is:

NPVbest = –$1675000 + 964176.855 * PVIFA

=–$1675000 + 964176.855 * 4.230538

= 2403986.82

OCFworst case = [(P – VC)Q – FC](1 –t) + Depreciation(t)

OCFWorst = [($32.355 – 23.54)(81900) – $852500](0.65) + 0.35($1675000/6) = $-84858.475 + 97708.33

OCF Worst = $ 12849.855

Now we can calculate the NPV using our Worst-case projections. There is no salvage value or NWC, so theNPV is:

NPV Worst = –$1675000 + 12849.855 * PVIFA

=–$1675000 + 12849.855 * 4.230538

= -1620638.2